OpenStudy (anonymous):

change to vertex form:FRACTION y= 1/5 (2x^2 -4x +7 )

OpenStudy (solomonzelman):

Posting from the previous question: y = ⅕(2x²-4x+7) at first I would factor out of 2. y = ⅖(x²-2x+3½) And now, you have inside the parenthesis 2½ more than what you want to make the parenthesis a perfect square Follow this carefully. y = ⅖(x²-2x+1+2½) y = ⅖(x²-2x+1)+(2½×⅖) y = ⅖(x²-2x+1)+({5/2}×{2/5}) y = ⅖(x²-2x+1)+(1) y = ⅖(x²-2x+1)+1 y = ⅖(x-1)²+1

OpenStudy (solomonzelman):

opens up. stretched horizontally. minimum point at (1,1)